OpenStudy (anonymous):

A quarter,a nickel, and a penny are flipped. Find he probability of each of the following. a.) The quarter shows head b.)The penny and nickel show heads c.)One coin shows heads d.)All three coins land te same way

OpenStudy (anonymous):

a) It's a coin flip! The probability is 1/2. b) Two coin flips, each one with a 1/2 chance of showing heads. The product is 1/4. c) This is trickier. There are 2*2*2 = 8 total possible outcomes. There are 3 such outcomes with 1 coin showing heads (HTT, THT, TTH). So, the probability is 3/8. d) See if you can do this one on your own, using the same method as for part c.

OpenStudy (perl):

d) HHH or TTT , which is 1/8 + 1/8 = 2/8